# An ordering for the strength of functional dependence

### v1

@@ Appendix B Proofs
(i): Due to sufficiency of $g$ for $Y|\mathbb{X}$ and of $h$ for $Y^{\prime}|\mathbb{X}^{\prime}\,$, we obtain from self-equitability of $\preccurlyeq_{ccx}$ (see Remark 1.5(d)) that $(Y,\mathbb{X})=_{ccx}(Y,g(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},h(\mathbb{X}^{\prime}))=_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,.$
(ii): Since $Y\uparrow_{st}g(\mathbb{X})$ and $Y^{\prime}\uparrow_{st}h(\mathbb{X}^{\prime})\,$, the (on $\mathsf{Ran}(F_{Y})\times\mathsf{Ran}(F_{g(\mathbb{X})})$ resp. $\mathsf{Ran}(F_{Y^{\prime}})\times\mathsf{Ran}(F_{h(\mathbb{X})})$ uniquely determined) copulas $C_{Y,g(\mathbb{X})}$ and $C_{Y^{\prime},h(\mathbb{X}^{\prime})}$ can chosen to be SI on $0,1]^{2}$. Hence, by Theorem 1.6, $C_{Y,g(\mathbb{X})}\leq_{c}C_{Y^{\prime},h(\mathbb{X}^{\prime})}$ and $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}}})$ imply $(Y,g(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},h(\mathbb{X}^{\prime}))\,.$ Then, the statement follows from part (i).
Statement (iii) follows from (ii) replacing $g(\mathbb{X})$ and $h(\mathbb{X}^{\prime})$ by $g(\mathbb{X})$ and $h(\mathbb{X}^{\prime})\,. \blacksquare$
@@
$P(Y\leq y\mid Z=q_{Z}(t))=P(Z+\sigma\varepsilon\leq y\mid Z=q_{Z}(t))=F_{\varepsilon}(\tfrac{y-q_{Z}(t)}{\sigma})\,$, (50)
@@
$t$ $\leq F_{Z}(\tfrac{\sigma y^{\prime}-\sigma^{\prime}y}{\sigma-\sigma^{\prime}})$ (51)
$\Longleftrightarrow$ $q_{Z}(t)$ $\leq\frac{\sigma y^{\prime}-\sigma^{\prime}y}{\sigma-\sigma^{\prime}}$ (52)
$\Longleftrightarrow$ $\frac{y-q_{Z}(t)}{\sigma}$ $\geq\frac{y^{\prime}-q_{Z}(t)}{\sigma^{\prime}}$ (53)
$\Longrightarrow$ $F_{\varepsilon}(\tfrac{y-q_{Z}(t)}{\sigma})$ $\geq F_{\varepsilon}(\tfrac{y^{\prime}-q_{Z}(t)}{\sigma^{\prime}}),$ (54)
@@
$\int_{0}^{x}P(Y\leq q_{Y}(v)\mid Z=q_{Z}(t))\mathsf{\,d}t\geq\int_{0}^{x}P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid Z=q_{Z}(t))\mathsf{\,d}t\quad\text{for all }x\in[0,1]$ (55)
with equality for x=1.x=1\,. Using Theorem 2.2, it follows that $(Y,f(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},f(\mathbb{X}))\,.$ Since $Y\mid f(\mathbb{X})$ is sufficient for $Y\mid\mathbb{X}$ and, similarly, $Y^{\prime}\mid f(\mathbb{X})$ is sufficient for $Y^{\prime}\mid\mathbb{X}$, we obtain from Theorem 3.3(i) that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$. $\blacksquare$
@@ B.4 Remaining proofs of Section 2
The equivalence of (i) and (ii) is explained in detail in Example 2.4 and Figure 1. The condition $q=q^{\prime}$ thereby equals the marginal constraint.
(iii): $X$ and $Y$ are independent if and only if $p_{00}=(1-p)(1-q)$ if and only if $p_{01}=p(1-q)$ if and only if $\alpha=\beta$.
(iv): $Y$ perfectly depends on $X$ if and only if (1) $p_{01}=0$, $p=q$, and $p_{00}=1-p$ or (2) $p_{00}=0$, p=1qp=1-q, and $p_{01}=p$ if and only if $\alpha\wedge\beta=0$ and $\alpha\vee\beta=1$.
(v): $X$ and $Y$ are comonotone if and only if $p_{00}=(1-p)\wedge(1-q)$ if and only if $\alpha=\frac{1-q}{1-p}\wedge 1$.
(vi): $X$ and $Y$ are countermonotone if and only if $p_{00}=((1-p)+(1-q)-1)\vee 0$ if and only if $\alpha=0\vee\left(1-\frac{q}{1-p}\right)$. $\blacksquare$
@@ B.5 Remaining proofs of Section 3
$(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\quad\Longleftrightarrow\quad(F_{Y}(Y),\Phi(S))\leq_{c}(F_{Y^{\prime}}(Y^{\prime}),\Phi(S^{\prime})).$ (56)
Since $(Y,S)$ is bivariate normal with $\mathrm{Var}(S)=1$, the distribution function of the random vector $(F_{Y}(Y),\Phi(S))$ is the Gaussian copula with correlation parameter $\varrho=\sqrt{\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}/\sigma_{Y}^{2}}\,;$ see (14). Similarly, $(F_{Y^{\prime}}(Y^{\prime}),\Phi(S^{\prime}))$ follows a Gaussian copula with parameter $\varrho^{\prime}=\sqrt{\Sigma_{Y^{\prime},\mathbb{X}^{\prime}}\Sigma_{\mathbb{X}^{\prime}}^{-}\Sigma_{\mathbb{X}^{\prime},Y^{\prime}}/\sigma_{Y^{\prime}}^{2}}$. Since the Gaussian copula is increasing in its parameter with respect to the concordance order [3], the right-hand side of (56) is equivalent to $\varrho\leq\varrho^{\prime}\,.$ $\blacksquare$
@@ B.6 Proofs of Section 4
To prove assertion (i), we use that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies (25) as already shown. Since the conditional convex order satisfies Axiom 3, $(Y,\mathbb{X})$ is a minimal element in $\preccurlyeq_{ccx}$ if $Y$ and $\mathbb{X}$ are independent. In this case, we have $\xi_{\varphi}(Y,\mathbb{X})=0=\Lambda_{\varphi}(Y,\mathbb{X})$. Similarly, the conditional convex order satisfies Axiom 4, and thus $(Y,\mathbb{X})$ is a maximal element in $\preccurlyeq_{ccx}$ if $Y$ perfectly depends on $\mathbb{X}$. In this case, we have $\xi_{\varphi}(Y,\mathbb{X})=1=\Lambda_{\varphi}(Y,\mathbb{X})$, which is also immediate from the definition of the functionals.
The zero-independence property in (ii) follows from the fact that $Y$ and $\mathbb{X}$ are independent if and only if $F_{Y|\mathbb{X}=\mathbb{x}}=F_{Y}$ for $P^{\mathbb{X}}$-almost every $\mathbb{x}\in\mathbb{R}^{p}$. This condition is equivalent to requiring that the inner integrals in the definitions of $\xi_{\varphi}$ and $\Lambda_{\varphi}$ vanish $P^{\mathbb{X}}$-almost everywhere and $P^{\mathbb{X}}\otimes P^{\mathbb{X}}$-almost everywhere, respectively, since $\varphi$ is strictly convex at 0 with $\varphi(0)=0$.
The max-functionality property of $\Lambda_{\varphi}$ is a direct extension of [5, Proof of Theorem 3.2] to a multivariate vector $\mathbb{X}=(X_{1},\ldots,X_{p})$. To prove max-functionality for $\xi_{\varphi}$, it remains to show that $\xi_{\varphi}(Y,\mathbb{X})=1$ implies $Y=f(\mathbb{X})$ almost surely. Therefore, define $f_{y}(\mathbb{x}):=F_{Y|\mathbb{X}=\mathbb{x}}(y)$ and $a:=F_{Y}(y)$ and set
@@
For verifying (iii), it remains to show that $\Lambda_{\varphi}(Y,\mathbb{X})=\Lambda_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ implies conditional independence of $Y$ and $\mathbb{Z}$ given $\mathbb{X}$. Therefore, define again $f_{y}(\mathbb{x}):=F_{Y|\mathbb{X}=\mathbb{x}}(y)$ and $g_{y}(\mathbb{x},\mathbb{z}):=F_{Y|\mathbb{X}=\mathbb{x},\mathbb{Z}=\mathbb{z}}(y)\,.$ Using $\Lambda_{\varphi}(Y,\mathbb{X})=\Lambda_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ and Jensen’s inequality, it follows that
$\beta_{\varphi}\Lambda_{\varphi}(Y,\mathbb{X})=\int\int\varphi(f_{y}(\mathbb{x}_{1})-f_{y}(\mathbb{x}_{2}))\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x}_{1})\otimes P^{\mathbb{X}}(\mathbb{x}_{2})\mathsf{\,d}P^{Y}(y)$
$=\int\int\varphi\left(\int g_{y}(\mathbb{x}_{1},\mathbb{z}_{1})-g_{y}(\mathbb{x}_{2},\mathbb{z}_{2})\mathsf{\,d}P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{1}}(\mathbb{z}_{1})\otimes P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{2}}(\mathbb{z}_{2})\right)\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x}_{1})\otimes P^{\mathbb{X}}(\mathbb{x}_{2})\mathsf{\,d}P^{Y}(y)$
$\leq\int\int\int\varphi(g_{y}(\mathbb{x}_{1},\mathbb{z}_{1})-g_{y}(\mathbb{x}_{2},\mathbb{z}_{2}))\mathsf{\,d}P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{1}}(\mathbb{z}_{1})\otimes P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{2}}(\mathbb{z}_{2})\mathsf{\,d}P^{\mathbb{X}}(\mathbb{x}_{1})\otimes P^{\mathbb{X}}(\mathbb{x}_{2})\mathsf{\,d}P^{Y}(y)$
@@
$\varphi(f_{y}(\mathbb{x}_{1})-f_{y}(\mathbb{x}_{2}))=\mathbb{E}\varphi\left(g_{y}(\mathbb{x}_{1},\mathbb{Z}_{\mathbb{x}_{1}}^{1})-g_{y}(\mathbb{x}_{2},\mathbb{Z}_{\mathbb{x}_{2}}^{2})\right)$ (57)
@@
$f_{y}(\mathbb{x}_{1})-g_{y}(\mathbb{x}_{1},\mathbb{Z}_{\mathbb{x}_{1}}^{1})=f_{y}(\mathbb{x}_{2})-g_{y}(\mathbb{x}_{2},\mathbb{Z}_{\mathbb{x}_{2}}^{2})\quad\text{a.s.}$ (58)
for all $(\mathbb{x}_{1},\mathbb{x}_{2})\in G_{y}$ and $y\in N^{c}\,.$ Integrating out in (58) then gives

### v2

@@ B Proofs
i: Due to sufficiency of $g$ for $Y|\mathbb{X}$ and of $h$ for $Y^{\prime}|\mathbb{X}^{\prime}\,$, we obtain from self-equitability of $\preccurlyeq_{ccx}$ (see Remark 1.5d) that $(Y,\mathbb{X})=_{ccx}(Y,g(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},h(\mathbb{X}^{\prime}))=_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\,.$
ii: Since $Y\uparrow_{st}g(\mathbb{X})$ and $Y^{\prime}\uparrow_{st}h(\mathbb{X}^{\prime})\,$, the (on $\mathsf{Ran}(F_{Y})\times\mathsf{Ran}(F_{g(\mathbb{X})})$ resp. $\mathsf{Ran}(F_{Y^{\prime}})\times\mathsf{Ran}(F_{h(\mathbb{X})})$ uniquely determined) copulas $C_{Y,g(\mathbb{X})}$ and $C_{Y^{\prime},h(\mathbb{X}^{\prime})}$ can chosen to be SI on $0,1]^{2}$. Hence, by Theorem 1.6, $C_{Y,g(\mathbb{X})}\leq_{c}C_{Y^{\prime},h(\mathbb{X}^{\prime})}$ and $\overline{\mathsf{Ran}(F_{Y})}=\overline{\mathsf{Ran}(F_{Y^{\prime}}})$ imply $(Y,g(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},h(\mathbb{X}^{\prime}))\,.$ Then, the statement follows from part i.
Statement iii follows from ii replacing $g(\mathbb{X})$ and $h(\mathbb{X}^{\prime})$ by $g(\mathbb{X})$ and $h(\mathbb{X}^{\prime})\,. \blacksquare$
@@
$P(Y\leq y\mid Z=q_{Z}(t))=P(Z+\sigma\varepsilon\leq y\mid Z=q_{Z}(t))=F_{\varepsilon}(\tfrac{y-q_{Z}(t)}{\sigma})\,$, (B.18)
@@
$t$ $\leq F_{Z}(\tfrac{\sigma y^{\prime}-\sigma^{\prime}y}{\sigma-\sigma^{\prime}})$ (B.19)
$\Longleftrightarrow$ $q_{Z}(t)$ $\leq\frac{\sigma y^{\prime}-\sigma^{\prime}y}{\sigma-\sigma^{\prime}}$ (B.20)
$\Longleftrightarrow$ $\frac{y-q_{Z}(t)}{\sigma}$ $\geq\frac{y^{\prime}-q_{Z}(t)}{\sigma^{\prime}}$ (B.21)
$\Longrightarrow$ $F_{\varepsilon}(\tfrac{y-q_{Z}(t)}{\sigma})$ $\geq F_{\varepsilon}(\tfrac{y^{\prime}-q_{Z}(t)}{\sigma^{\prime}}),$ (B.22)
@@
$\int_{0}^{x}P(Y\leq q_{Y}(v)\mid Z=q_{Z}(t))\mathsf{\,d}t\geq\int_{0}^{x}P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid Z=q_{Z}(t))\mathsf{\,d}t\quad\text{for all }x\in[0,1]$ (B.23)
with equality for x=1.x=1\,. Using Theorem 2.2, it follows that $(Y,f(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},f(\mathbb{X}))\,.$ Since $Y\mid f(\mathbb{X})$ is sufficient for $Y\mid\mathbb{X}$ and, similarly, $Y^{\prime}\mid f(\mathbb{X})$ is sufficient for $Y^{\prime}\mid\mathbb{X}$, we obtain from Theorem 3.3i that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$. $\blacksquare$
@@ B.4 Remaining proofs of Section 2
The equivalence of i and ii is explained in detail in Example 2.4 and Figure 1. The condition $q=q^{\prime}$ thereby equals the marginal constraint.
iii: $X$ and $Y$ are independent if and only if $p_{00}=(1-p)(1-q)$ if and only if $p_{01}=p(1-q)$ if and only if $\alpha=\beta$.
iv: $Y$ perfectly depends on $X$ if and only if (1) $p_{01}=0$, $p=q$, and $p_{00}=1-p$ or (2) $p_{00}=0$, p=1qp=1-q, and $p_{01}=p$ if and only if $\alpha\wedge\beta=0$ and $\alpha\vee\beta=1$.
v: $X$ and $Y$ are comonotone if and only if $p_{00}=(1-p)\wedge(1-q)$ if and only if $\alpha=\frac{1-q}{1-p}\wedge 1$.
vi: $X$ and $Y$ are countermonotone if and only if $p_{00}=((1-p)+(1-q)-1)\vee 0$ if and only if $\alpha=0\vee\left(1-\frac{q}{1-p}\right)$. $\blacksquare$
@@ B.5 Remaining proofs of Section 3
$(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\quad\Longleftrightarrow\quad(F_{Y}(Y),\Phi(S))\leq_{c}(F_{Y^{\prime}}(Y^{\prime}),\Phi(S^{\prime})).$ (B.24)
Since $(Y,S)$ is bivariate normal with $\mathrm{Var}(S)=1$, the distribution function of the random vector $(F_{Y}(Y),\Phi(S))$ is the Gaussian copula with correlation parameter $\varrho=\sqrt{\Sigma_{Y,\mathbb{X}}\Sigma_{\mathbb{X}}^{-}\Sigma_{\mathbb{X},Y}/\sigma_{Y}^{2}}\,;$ see (17). Similarly, $(F_{Y^{\prime}}(Y^{\prime}),\Phi(S^{\prime}))$ follows a Gaussian copula with parameter $\varrho^{\prime}=\sqrt{\Sigma_{Y^{\prime},\mathbb{X}^{\prime}}\Sigma_{\mathbb{X}^{\prime}}^{-}\Sigma_{\mathbb{X}^{\prime},Y^{\prime}}/\sigma_{Y^{\prime}}^{2}}$. Since the Gaussian copula is increasing in its parameter with respect to the concordance order [5], the right-hand side of (B.24) is equivalent to $\varrho\leq\varrho^{\prime}\,.$ $\blacksquare$
@@ B.6 Proofs of Section 4
To prove assertion i, we use that $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies (29) as already shown. Since the conditional convex order satisfies Axiom 3, $(Y,\mathbb{X})$ is a minimal element in $\preccurlyeq_{ccx}$ if $Y$ and $\mathbb{X}$ are independent. In this case, we have $\xi_{\varphi}(Y,\mathbb{X})=0=\Lambda_{\varphi}(Y,\mathbb{X})$. Similarly, the conditional convex order satisfies Axiom 4, and thus $(Y,\mathbb{X})$ is a maximal element in $\preccurlyeq_{ccx}$ if $Y$ perfectly depends on $\mathbb{X}$. In this case, we have $\xi_{\varphi}(Y,\mathbb{X})=1=\Lambda_{\varphi}(Y,\mathbb{X})$, which is also immediate from the definition of the functionals.
The zero-independence property in ii follows from the fact that $Y$ and $\mathbb{X}$ are independent if and only if $F_{Y|\mathbb{X}=\mathbb{x}}=F_{Y}$ for $P^{\mathbb{X}}$-almost every $\mathbb{x}\in\mathbb{R}^{p}$. This condition is equivalent to requiring that the inner integrals in the definitions of $\xi_{\varphi}$ and $\Lambda_{\varphi}$ vanish $P^{\mathbb{X}}$-almost everywhere and $P^{\mathbb{X}}\otimes P^{\mathbb{X}}$-almost everywhere, respectively, since $\varphi$ is strictly convex at 0 with $\varphi(0)=0$.
The max-functionality property of $\Lambda_{\varphi}$ is a direct extension of [4, Proof of Theorem 3.2] to a multivariate vector $\mathbb{X}=(X_{1},\ldots,X_{p})$. To prove max-functionality for $\xi_{\varphi}$, it remains to show that $\xi_{\varphi}(Y,\mathbb{X})=1$ implies $Y=f(\mathbb{X})$ almost surely. Therefore, define $f_{y}(\mathbb{x}):=F_{Y|\mathbb{X}=\mathbb{x}}(y)$ and $a:=F_{Y}(y)$ and set
@@
For verifying iv, it remains to show that $\Lambda_{\varphi}(Y,\mathbb{X})=\Lambda_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ implies conditional independence of $Y$ and $\mathbb{Z}$ given $\mathbb{X}$. Therefore, define again $f_{y}(\mathbb{x}):=F_{Y|\mathbb{X}=\mathbb{x}}(y)$ and $g_{y}(\mathbb{x},\mathbb{z}):=F_{Y|\mathbb{X}=\mathbb{x},\mathbb{Z}=\mathbb{z}}(y)\,.$ Using $\Lambda_{\varphi}(Y,\mathbb{X})=\Lambda_{\varphi}(Y,(\mathbb{X},\mathbb{Z}))$ and Jensen’s inequality, it follows that
$\beta_{\varphi}\Lambda_{\varphi}(Y,\mathbb{X})=\int\int\varphi(f_{y}(\mathbb{x}_{1})-f_{y}(\mathbb{x}_{2}))\mathsf{\,d}P^{\mathbb{X}}\otimes P^{\mathbb{X}}(\mathbb{x}_{1},\mathbb{x}_{2})\mathsf{\,d}P^{Y}(y)$
$=\int\int\varphi\left(\int g_{y}(\mathbb{x}_{1},\mathbb{z}_{1})-g_{y}(\mathbb{x}_{2},\mathbb{z}_{2})\mathsf{\,d}P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{1}}\otimes P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{2}}(\mathbb{z}_{1},\mathbb{z}_{2})\right)\mathsf{\,d}P^{\mathbb{X}}\otimes P^{\mathbb{X}}(\mathbb{x}_{1},\mathbb{x}_{2})\mathsf{\,d}P^{Y}(y)$
$\leq\int\int\int\varphi(g_{y}(\mathbb{x}_{1},\mathbb{z}_{1})-g_{y}(\mathbb{x}_{2},\mathbb{z}_{2}))\mathsf{\,d}P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{1}}\otimes P^{\mathbb{Z}|\mathbb{X}=\mathbb{x}_{2}}(\mathbb{z}_{1},\mathbb{z}_{2})\mathsf{\,d}P^{\mathbb{X}}\otimes P^{\mathbb{X}}(\mathbb{x}_{1},\mathbb{x}_{2})\mathsf{\,d}P^{Y}(y)$
@@
$\varphi(f_{y}(\mathbb{x}_{1})-f_{y}(\mathbb{x}_{2}))=\mathbb{E}\varphi\left(g_{y}(\mathbb{x}_{1},\mathbb{Z}_{\mathbb{x}_{1}}^{1})-g_{y}(\mathbb{x}_{2},\mathbb{Z}_{\mathbb{x}_{2}}^{2})\right)$ (B.25)
@@
$f_{y}(\mathbb{x}_{1})-g_{y}(\mathbb{x}_{1},\mathbb{Z}_{\mathbb{x}_{1}}^{1})=f_{y}(\mathbb{x}_{2})-g_{y}(\mathbb{x}_{2},\mathbb{Z}_{\mathbb{x}_{2}}^{2})\quad\text{a.s.}$ (B.26)
for all $(\mathbb{x}_{1},\mathbb{x}_{2})\in G_{y}$ and $y\in N^{c}\,.$ Integrating out in (B.26) then gives